LOG-IN
Displaying reports 1841-1860 of 3359.Go to page Start 89 90 91 92 93 94 95 96 97 End
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 21:14, Monday 04 November 2019 (1798)Get code to link to this report
Improved damping filter for PR pitch

A filter improvement (suggested by Matteo B.) has been implemented for PR pitch. I also added some notch at the frequency of the main lines.

Pic 1 comparison of the filters zpk

Pic 2 performances of the new filter

It seems to work very well.

Images attached to this report
1798_20191104131310_cfr.png 1798_20191104131347_newfilter.png
R&D (FilterCavity)
Print this report.
NaokiAritomi - 23:50, Friday 01 November 2019 (1797)Get code to link to this report
Locking accuracy with good IR alignment

First I checked IR alignment and found that IR TEM00 transmission was small. So I tweaked steering mirrors for IR injection and could recover the good alignment. Mode matching is around 95%.

Mode IR transmission
TEM00 385
pitch 110
IG20 -
offset 94

Then we measured locking accuracy again as entry 1769

Calibration:
peak-peak voltage of error signal: 266 mV
peak-peak time difference of error signal: 114 ms
slope = peak-peak voltage of error signal/peak-peak time difference of error signal = 266 mV/114 ms = 2.33 V/s
AOM scan frequency: 500 mHz
AOM deviation: 1 kHz
AOM scan speed for green: AOM deviation/AOM scan frequency = 2 kHz/s
AOM scan speed for IR: 2 kHz/s /2 = 1 kHz/s
 
Calibration factor:
AOM scan speed (Hz/s)/(2*slope(V/s)) = 1000(Hz/s)/(2*2.33(V/s)) = 215 Hz/V
 
Attached picture shows IR error signal. Locking accuracy is 4.4 Hz which corresponds to 4.7 pm. 1.3 Hz out of 4.4 Hz is coming from 600 Hz turbo pump and 1 Hz out of 4.4 Hz is coming from 9 Hz input mirror pitch (entry 1795). If we remove the 9 Hz and 600 Hz peak, we will have 2.1 Hz of locking accuracy and squeezing degradation will be much less.
Images attached to this report
1797_20210724224021_lockingaccuracy20191101.png
KAGRA MIR (Polarization)
Print this report.
PengboLi - 21:00, Friday 01 November 2019 (1796)Get code to link to this report
Polarization Maps on TAMA#1
[Simon, Pengbo]

We measured the Polarization map on TAMA#1 with different polarization angles. O degrees represents the pure P- polarization map.
As can be seen, these images show the same pattern in structure. Also the S- and P- polarization both show an apparent offset, which is most likely due to the birefringence effect.

Images attached to this report
1796_20191101125712_figurep1.png 1796_20191101125716_figures1.png 1796_20191106033706_figurep2.png 1796_20191106033716_figures2.png
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 15:36, Friday 01 November 2019 (1795)Get code to link to this report
Input mirror intermediate mass touching issue

[Aritomi, Matteo, Eleonora, Takahashi]

Yesterday we opened the input chamber with the help of Takahashi-san, and we confirm the intermediate mass was touching and picomotors were at the end range. (See entry #1783)

We adjusted IM mass and moved the picomotors by hand (and josystic) to make the reflection from the input mirror to superpose with the incoming one. It was quite hard to achive the superposition as we didn't notice that the intemediate mass was touching again and the picomotors couldn't move the mirror properly. Also the gatevalve between input and BS distorts the beam and causes multiple reflection.

Pic 1 and 2 show the current position of picomotors. In the pitch case we are quite close to the end of the range.

Today, since the vacuum was restored, we open the gatevalves and we could realign and lock the cavity again.

Anyway we found that the spectrum of the pitch motion has a large, sharp peak at 9 Hz that was not there before (pic3) . The TF seems fine (pic4). I wonder if there is still a problem with IM mass. As a first step I will try to adjust the control to damp it better.

Images attached to this report
1795_20191101073939_picojaw.jpg 1795_20191101073952_picopitch.jpg 1795_20191101132117_imspe.png 1795_20191101132129_imspetf.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 13:42, Friday 01 November 2019 (1794)Get code to link to this report
Strange sound from DDS board (recently happened after we just switch on DDS board)

Recently we switched off DDS board many times for the sake of avoiding RF signal cross-talk.

But recently every time we switch on again DDS, we heard a strange sound. From the sound, we guess the frequency seems to be a fixed audio frequency.

Up to now, our solution is

1. switch it off the rack containing the DDS board.

2. take out DDS3 board

3. switch on the rack

4. Put back DDS3 board while the power is on

By following this procedure, we could avoid the sound problem. But we still don't know what is the reason for this sound.

R&D (FilterCavity)
Print this report.
YaoChinHuang - 23:17, Thursday 31 October 2019 (1792)Get code to link to this report
Test of QPD by using 1064nm light

[Yao-Chin, Aritomi]

Testing method was same as entry 1785. However, we changed laser source to infrared light of 1064 nm. Infrared light hit in the first quarter of QPD2 (QPD2 input 1). We connected the corresponding DC output to the oscilloscope to check DC voltage. We also connected the corresponding RF output to a 32 dB amplifier and then check its power spectrum by a spectrum analyzer (Keysight N9320B). 

We measured the light power, DC voltage, DC current by using "PuTTY", and RF channel power spectrum with RBW of 1MHz and average 28 times shown in pic. 1. Dark noise of pic.1 included the 32 dB amplifier and instrument when no IR light hit to QPD. In this measurement, when light power goes up, power loss is going down unlike the measurement in entry 1785. Green power in entry 1785 was too much since maximum DC current is 10mA according to specification and calculated DC current is already more than 10mA with 53.5 mW of green.

Pic. 2 shows same measurement in NIKHEF. The result is similar to our result, but we have some peaks around 25MHz, 125MHz, 140MHz.

Light Power

[mW]

DC Voltage

[mV]

DC Current

[mA]

calculated DC current (mA)

(photosensitivity is 0.55A/W

according to specification)

power loss (%)

(ratio of measured DC current/calculated DC current)

0.58 248 0 0.319 100
1.25 504 0.1 0.688 85
2.5 1000 0.6 1.38 57
3.7 1500 1.1 2.04 46
4.5 1940 1.5 2.48 40
7.3 3000 2.7 4.02 33
9.5 4000 3.7 5.23 29
11.7 5000 4.8 6.44 25
13.5 5960 5.7 7.43 23
15.5 7040 6.8 8.53 20
17 8000 7.7 9.35 18
Images attached to this report
1792_20191031145702_aa.jpg
Non-image files attached to this report
KAGRA MIR (Absorption)
Print this report.
PengboLi - 16:55, Thursday 31 October 2019 (1789)Get code to link to this report
Reflectance and Transmittance of the OSTM sample with ATC coatings
[Simon, Pengbo]

Today we reconfigured the setup to birefringence measurement. First, we measured the reflectance and transmittance under different conditions.

With the P-pol input polarization, we got
P_in = 7.64 mW, P_refl = 7.559 mW, P_trans = 0.206 mW
R = 0.9894, T = 0.0270
With the S-pol input polarization, we got
P_in = 7.649 mW, P_refl = 7.602 mW, P_trans = 0.007 mW
R = 0.9939, T = 0.0009

Then, we did the S-polarization map on TAMA#1. The result will be shown tomorrow.

KAGRA MIR (Absorption)
Print this report.
PengboLi - 09:40, Thursday 31 October 2019 (1788)Get code to link to this report
Absorption map of OSTM with ATC coating
[Simon, Pengbo]

We reconfigured the PCI system and measured the absorption map of the OSTM sample with ATC coating.

As can be seen, the figure shows a large value of absorption. And there are many spots with absorption excesses. Most likely, they are due to defects of the coating
Images attached to this report
1788_20191031051107_figureabs1.png 1788_20191031051114_figureabs2.png
R&D (FilterCavity)
Print this report.
YaoChinHuang - 22:23, Wednesday 30 October 2019 (1786)Get code to link to this report
Measure demodulator performance

[Yao-Chin, Yuhang]

Frequency from DDS with -6.5dBm sent to input LO & RF ports of demodulator (ZFMIQ-70D). The demodulator output I & Q ports connected low pass filter of DC-2.5MHz. We measured the Vpp and phase difference of I & Q ports from oscilloscope. 

Note: Unbalance Amplitude (dB)= 20* log (Vpp of I port/ Vpp of Q port)

(I) Fixed LO port frequency of 78 MHz

LO 
(MHz)
RF
(MHz)
Vpp of I
(mV)
Vpp of Q
(mV)
Unbalance Amplitude
(dB)
Phase Difference btw I&Q
(degree)
78
78.0001
286
254
1.03
90.4
78
78.001
284
254
0.97
90.8
78
78.01
282
250
1.05
90.2
78
78.05
255
222
1.20
90.5
78
78.1
210
182
1.82
91.7

 
(II) Fixed difference frequency of 100 Hz between LO and RF ports tune range from 66 to 80 MHz.

LO
(MHz)
RF
(MHz)
Vpp of I
(mV)
Vpp of Q
(mV)
Unbalance Amplitude
(dB)
Phase Difference btw I&Q
(degree)
66.0001
66
260
260
0
91.2
67.0001
67
262
260
0.07
90.1
68.0001
68
266
260
0.20
90.9
69.0001
69
266
260
0.20
90.6
70.0001
70
270
260
0.33
90.7
71.0001
71
272
260
0.39
90.9
72.0001
72
274
260
0.46
90.9
73.0001
73
276
259
0.55
90.6
74.0001
74
278
258
0.65
89.8
75.0001
75
281
258
0.74
90.2
76.0001
76
282
258
0.77
90.2
77.0001
77
282
255
0.87
90.6
78.0001
78
286
254
1.03
90.4
79.0001
79
286
254
1.03
90.1
80.0001
80
286
251
1.13
90.9

* Phase: Q=I+90o for LO>RF, Q=I-90o for LO<RF

R&D (FilterCavity)
Print this report.
YuhangZhao - 21:39, Wednesday 30 October 2019 (1785)Get code to link to this report
Test of QPD by using 532nm light

To have an idea about how PD responses (at RF frequency) to green light. We did this test.

The test set-up is as shown in the attached figure 1 and 2. We put a periscope after the first green FI. It brings the green beam up to the second layer of the bench. At the second layer, we put a lens and a steering mirror to make sure the light is small enough and hit in the first quarter of QPD2. We connected the corresponding DC channel to the oscilloscope to check the output voltage DC. We also connect the corresponding RF channel to a 32dB amplifier and then a spectrum analyzer(Keysight N9320B) to check its noise spectrum.

To avoid any modulation and even RF cross-talk, we turned off the DDS system. The green power is changed by moving the offset of high voltage driver(connect to the PZT of SHG). We checked this offset is stable enough so that the green power doesn't change larger than 4% within the time of one measurement.

Within one measurement, we measured the light power/output voltage DC/output current(measured by using the 'putty')/RF channel spectrum.

Output voltage DC(V) Light power(mW) Output current(mA)

Calculated current(mA)

(photosensitivity is 0.2A/W according to specification)

power loss (%)

(inferred from measured and calculated current)

5.6 34.5 5.3 6.9 23.2
7.7 53.5 7.4 10.7 30.8
8.8 70 8.7 14 37.9
10 90 10 18 44

The relation between power and voltage/current is attached in figure 4.

I also calculated the shot noise by using this formula: shot noise = 10*log((sqrt(2*1.6e-19*I)*sqrt(RBW))**2*35/1mW)+32+22, then I got the calculated shot noise level

Output Voltage DC(V) Calculated shot noise(dBm)
5.6 -48.3
7.7 -46.8
8.8 -46.1
10 -45.5

While the measured shot noise is attached in figure 3, which shows -60dBm for all different light power.

Images attached to this report
1785_20191030131525_setup1.jpeg 1785_20191030131530_setup2.jpeg 1785_20191030131715_good2.png 1785_20191030133809_figure3.png
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 19:44, Wednesday 30 October 2019 (1783)Get code to link to this report
Input picomotor stuck
Today when we tried to realign the cavity we found that we have to steer a lot the input mirror in yaw.  Maybe it moved a bit during the opening/closing of the chamber.
 
We tried to move the mirror with picomotors but they got stuck and now the intermediate mass is touching the frame, as it can be seen from the spectrum attached.
 
Probabily we need to open the chamber to fix it.
Images attached to this report
1783_20191030114419_input.png
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 19:37, Wednesday 30 October 2019 (1782)Get code to link to this report
Top coil of input mirror fixed

On Tue 29/11 we open INPUT vacuum chamber. Takahashi-san found the wire to the ground pin of the circuit of the top coil was broken. See picture.

He repaired it. We closed the chamber and check the the coil is working fine now.

Images attached to this report
1782_20191030113743_in4.jpeg
R&D (FilterCavity)
Print this report.
YuhangZhao - 17:16, Wednesday 30 October 2019 (1780)Get code to link to this report
Lens position before homodyne

[Yuhang, Aritomi]

We had homodyne noise spectrum bump problem reported in entry1529.

One of the reasons could be the cut issue of PD, and I checked the entry1159. I found that the 'beam analyzer' was put in a wrong position. The corrected version is attached here. In this version, the 'beam analyzer' is marking the position where the beam size(100um) is 5times smaller than the size of PD(500um). So we can decide that the homodyne PD should be 26mm~32mm after 30mm lens. This means we have only 6mm range to make sure the beam is smaller enough than the size of PD(500um).

We also measured again the shot noise after this work(attached picture 2 and 3). The bump issue sometimes is till present, but it covers region smaller than 20Hz and rarely show up. But we found small 45Hz and 54Hz peaks which are not present in previous measurement. They may due to we didn't cover bench as before(now the west side is half open and south side is open).

Images attached to this report
1780_20191030090244_20191030homosize.png 1780_20191030090611_celingon.jpg 1780_20191030090614_off.jpg
R&D (FilterCavity)
Print this report.
MatteoLeonardi - 17:03, Wednesday 30 October 2019 (1781)Get code to link to this report
Second failure of the BS TMP
[Eleonora, Takahashi, Matteo]

Today we had another failure of the BS TMP.
The error is number 089 which is for "Rotor vibration". The pump stopped autonomously.

After consulting with Takahashi-san we closed the gate valve between the TMP and the BS chamber as well as the valve between the BS TMP and RP and switched off the BS RP.
At this stage the pressure in the BS+PR was around 10^-3 mbar. In the input tower it was 10^-7 and in the arm 10^-8 (the gate valve around the input tower were still closed due to the work performed yesterday on the input mirror coil).
We open the gate valve between input and PB+PR and the pressure stabilized around 10^-4. So we let it go down few minutes until the pressure in input+BS+PR reached 10^-6 and then we opened the gate valve between the arm and the CITF. After that the pressure stabilized around 10^-6 and it is slowly going down.

In this moment there is only the pumping system of the input tower working for the central area.
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 14:08, Wednesday 30 October 2019 (1779)Get code to link to this report
Comment to Work on Cryostat (Click here to view original report: 1768)

This entry is log on 28th Oct.
I cut some teflon sheet for cushioning between glass window and adapter.

Images attached to this comment
1779_20191030060427_20191029teflon.jpg
KAGRA MIR (Polarization)
Print this report.
PengboLi - 11:21, Wednesday 30 October 2019 (1778)Get code to link to this report
Polarization Maps on OSTM with ATC coatings
[Simon, pengbo]

We measured the polarization map of OSTM with different polarization angles.
0 degrees represent pure p-polarization. 90 degrees represent pure s-polarization. 40.5 degrees represent a mixture of sp polarization with a ratio of 1 to 1.

Images attached to this report
1778_20191030032021_figurep1.png 1778_20191030032026_figures1.png 1778_20191101064337_figuresp1.png 1778_20191101064407_figurep2.png 1778_20191101064412_figures2.png 1778_20191101064417_figuresp2.png
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 10:38, Wednesday 30 October 2019 (1777)Get code to link to this report
FDS measurement: fit for different HOM angles

[Matteo, Eleonora]

We used the code that simulates the homodyne spectrum for FDS at differente angles (entry #1774) and tried to optimize the angle to fit the data of our last FDS measurement (entry #1751)

The results in the attached figure show a quite good agreement between data (solid line) and simulation (dotted line)  

The degradation parameters used in the code are the following:

sqz_dB = 16;                         % produced SQZ

L_rt = 100e-6;                       % FC losses

L_inj = 0.33;                        % Injection losses

L_ro = 0.11;                         % Readout losses

A0 = 0.1;                            % Squeezed field/filter cavity mode mismatch losses

C0 = 0.05;                           % Squeezed field/local oscillator mode mismatch losses

ERR_L =   5e-12;                     % Lock accuracy [m]

ERR_csi = 80e-3;                     % Phase noise[rad]

det =  50.62e3;                      % detuning frequency 

int = 2e3;                           % frequency range = det+/-int   

t_in_q   = 0.0014;                   % input mirror transmission 

gamma_fc = ((t_in_q + L_rt)/2)*fsr;  % = 59.6*2*pi

Images attached to this report
1777_20191030023841_fdsplotsim.png
R&D (FilterCavity)
Print this report.
NaokiAritomi - 22:22, Monday 28 October 2019 (1776)Get code to link to this report
Effect of second harmonic of 7MHz (CCFC)

To check how much second harmonic of 7MHz CC/LO beat note affects 14MHz CCSB beat note, I put TAMA RFPD before AMC and detected LO and CC at the same time. CC is directly injected from OPO and CC1 is locked. I compared 14MHz peak height with only CC or CC+LO. Pic. 1,2 shows 14MHz signal with only CC or CC+LO. Apparently 14MHz peak height is almost same.

Then I demodulated the 14MHz signal with 14MHz and measured spectrum of the demodulated signal (Pic. 3). Spectrum with CC+LO has some excess noise compared with only CC and the noise shape seems similar to CC2 phase noise. This noise should come from second harmonic of 7MHz.

Images attached to this report
1776_20191028142229_cconly.jpg 1776_20191028142235_cclo.jpg 1776_20191029071616_cclo20191028.png
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 18:12, Monday 28 October 2019 (1775)Get code to link to this report
Comment to Work on Cryostat (Click here to view original report: 1768)

Today I temporary installed the adapter flange and 4K shiled inside the cryostat chamber to confirm the screw hole position is correct.
The adapter flange and shield seemed to be O.K., but I need 10 more ultrasonic cleaned M6 screws.

In addition to that, we need super insulator on 4K shield and tape to fix cables.

The procudure for remaining tasks are:

  1. Prepare windows for 80K shield
  2. Install windows on 80K shield
  3. Cabling on 4K shield
  4. Attach SI on 4K shield
  5. Install 4K shield
Images attached to this comment
1775_20191028101038_20191028adapter.jpg 1775_20191028101042_20191028shield.jpg
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 15:35, Monday 28 October 2019 (1774)Get code to link to this report
Simulation of the HOM output for different detection angles

I adapted the code to simulate the squeezing degradation in the case of standalone filter cavity.  Basically, I removed the transfer matrix accountig for ITF  in the quantum noise computation.

Now if we select a homodyne detection angle the code should simulated what we exepected to see when taking the spectrum of the Homodyne output.

In the code I also removed the approximation of small detuning and I simulate a large detunig case (50kHz) as the one we tested recently. (Entries:  #1747#1751)

Plot 1 shows the expected results for Homodyne angle: 0 deg and 90 deg. Plot 2 shows an intermediate case (30 deg) compared to the  0 deg case.

The degradation parameters are the same reported in entry  #1766

Images attached to this report
1774_20191028073149_hom090.png 1774_20191028073326_hom030.png